OpenStudy (caerus):

TWO LOADS ARE CONNECTED IN PARALLEL TO AN A-C SOURCE AND TAKE CURRENTS OF 10 AND 40 AMP, RESPECTIVELY.IF THE SINUSOIDAL WAVE OF THE COMPONENT CURRENTS ARE OUT OF PHASE BY 30 ELEC DEG. CALCULATE THE RESULTANT CURRENT.

OpenStudy (caerus):

@CGGURUMANJUNATH @3mar @IrishBoy123

OpenStudy (caerus):

@inkyvoyd @sweetburger

OpenStudy (caerus):

ans= 48.7 A base on the book

OpenStudy (osprey):

@Caerus 10 cos 30 + 40 = 48.7 is my first stab at this one ..

OpenStudy (irishboy123):

I'm pretty sure that book answer is wrong. I threw the kitchen sink at this earlier and now I think @osprey has identified why there is an error :) Think of it, even if there was a rationale for adding the current vectors this way to get the **right** answer, how do you choose between: \(10 \cos 30 + 40\) and \(40 \cos 30 + 10\) and \(40 \sin 30 + 10\) etc. Where is the extra asymmetry or information that pinpoints one of these approaches as the correct over the other one**s**?! The easiest way to add these vectors is to first draw a phasor diagram. Once you have built a phasor, here we add them as follows: \(|I| = \sqrt { (40 \cos 30 + 10)^2 + (40 \sin 30)^2 } \approx 48.9 A\)

OpenStudy (osprey):

@IrishBoy123 That was my concern too ... which to "recommend" ?